every time the pedals go through a 360 degree rotation on a certain bike, the tires rotate 3 times. if the tires are 24 inches in diameter, what is the minimum number of complete rotations of the pedals needed for the bike to travel atleast 1 mile?
1) 12
2) 281
3) 561
4) 5,280

Answer by mananth(16949)

The diameter of the tyre is 24 inches
radius = 12 inches.
the circumference = 2*pi*r
=2*pi*12
=75.4 inches
If the cycle has to go 1 mile it has to go 63360 inches
the number of revolutions the tyre has to make = 63360 / 75.4
= 840.31 revolutions
if the tyre makes 840.31 revolutions the the pedal has to make
840.31 / 3
= 280.1 revolutions
OR 281 revolutions